# Initializing Services in Interactive ML Systems for Diverse Users

### v1

@@ 5 Experiments
where $r_{i}$ is the true user ratings and $\Omega_{i}$ is the list of movies rated by the user for subpopulation $i$. User dissatisfaction for a recommendation model, parameterized by $V\in\mathbb{R}^{d\times m}$, is captured by the excess error, namely
$\mathcal{L}_{i}(V,V_{i})=\|(U_{i}V)_{\Omega_{i}}-r_{i}\|_{2}^{2}-\|(U_{i}V_{i})_{\Omega_{i}}-r_{i}\|_{2}^{2}.$ (4)
We chose $d=5$ for the experiment. As described for the census experiment, Algorithm 1 iteratively selects subpopulations with probability proportional to their current excess error, and then solves the matrix factorization problem for the selected subpopulation, and adds the learnt item embedding to the list of services.
As shown in Figure 3, the total excess error achieved by Algorithm 1 is significantly lower than uniformly sampling users for their preferences. It is interesting to note that to achieve the same total excess error for five services initialized by our algorithm, uniform data collection and initialization would need 30 services.
6 Related Work
Retention.
User retention in machine learning systems is closely related to the decision dynamics between the provider and users studied in the single service setting by Hashimoto et al. (2018) and Zhang et al. (2019), and multiple services setting by Dean et al. (2022). In settings with multiple sub-populations of users of different types, the question of retention has been explored in parallel with the issue of fairness. The final outcome of such dynamics are heavily influenced by the initial data available to provider and the initial configuration (or partitioning) of the users across the offered services. Our work addresses these issue with theoretical guarantees.
Multiple Model Specialization.
Our work also considers the inherent heterogeneity in user preferences, especially prominent in distributed learning settings where data sources are users’ personal devices. For applications such as recommendation systems and personalized advertising, the utilization of multiple specialized models can yield improved predictions and outcomes. Users are often grouped into clusters, each representing unique interests and preferences. For instance, in movie recommendation systems, these clusters could represent users interested in different movie genres, or different combination of features. This approach has been adopted in recent federated learning formulations such as clustered federated learning (Sattler et al., 2020; Mansour et al., 2020) and is integral to the federated learning model of Ghosh et al. (2020), where users choose models for which to provide updates. This model specialization is also explored in the mixture of experts literature (see, e.g., Masoudnia and Ebrahimpour (2014)), which uses multiple ‘expert’ models to enhance accuracy and robustness, assigning inputs to the most suitable expert based on their features through a gating mechanism.
Clustering.
The $k$-means clustering problem is one of the most commonly encountered unsupervised learning problems. Given a set of known $n$ points in Euclidean space, the goal is to partition them into $k$ clusters (each characterized by a center), such that the sum of square of distances to their closest center is minimized. Dasgupta (2008) and Aloise et al. (2009) showed that the $k$-means problem is NP-Hard.
The most popular heuristic for $k$-means is Lloyd’s algorithm (Lloyd, 1982), often referred to as the "$k$-means algorithm", which proceeds by randomly initializing $k$ centers and then uses iterative updates to find a locally optimal $k$-means clustering, which can be arbitrarily bad compared to the globally optimal clustering. The performance of the $k$-means algorithm relies crucially on the initialization. Arthur and Vassilvitskii (2007) and Ostrovsky et al. (2013) proposed an elegant polynomial time algorithm for initializing centers, known as $k$-means++. Arthur and Vassilvitskii (2007) proved that the expected cost of the initial clustering obtained by $k$-means++ is at most $8(2+\log k)$ times the cost of optimal $k$-means clustering and showed a family of $k$-means instances where the approximation ratio is $2\log k$ showing their analysis is almost tight. Makarychev et al. (2020) improved the analysis to show an approximation ratio $5(2+\log k)$ and showed a family of instances with $5\log k$ approximation ratio, thus showing $k$-means++ is tight.
Our work generalizes the analysis of Arthur and Vassilvitskii (2007) to the setting where user’s preferences are represented as unknown points and the loss functions are unknown and not necessarily identical as long they satisfy Assumptions 1 and 2. Recent works in clustering (Kanungo et al., 2002; Ahmadian et al., 2019; Lattanzi and Sohler, 2019) provide a constant approximation ratio, and although they are polynomial time algorithms, these methods are data inefficient and rely on knowing the points a priori. Finally, another common clustering algorithm is Gaussian mixture models (GMMs), where each cluster is represented by a Gaussian distribution, thus clusters themselves can be viewed as ellipsoidal (which is a different goal from ours even we choose Mahalanobis distances as losses, for us in this case every point (user) is associated with a different loss geometry).
Facility Location Problem.
Our algorithm has some resemblances with the facility location problem where there are $n$ users, and a provider can set up at most $k$ facilities at one of $m$ candidate locations. This is also often known as the $k$-medoids objective (Arora et al., 2016). One of the key differences is the provider can choose from $m$ predecided locations, compared to our setting where the optimization space is infinite. However our algorithm initializes only at $k$ of the $n$ user preferences, hence it can be viewed as the candidate locations simply being the user preferences and thus $m=n$. A typical greedy algorithm for the $k$-medoids objective proceeds by evaluating the marginal decrement in total loss for all possible candidates and selects the candidate with maximal loss reduction. Thus (i) the algorithm would need to know all $\{\phi_{1},\ldots,\phi_{n}\}$ apriori, and (ii) deploy $n$ services to obtain all $n^{2}$ function evaluations $\mathcal{L}_{j}(\phi_{i},\phi_{j}),\;\forall i,j\in[n]$. This is infeasible in applications like online recommendation systems where $n$ is very large and typically a provider has the capacity to deploy only $k\ll n$ services. Hence a $k$-medoids like objective will not be reasonable in our setup with incomplete information.
Bose et al. (2022) studied the case where the provider had prior access to only $N<n$ users’ utility functions before deploying services and the goal was to minimize the worst case total (over all $n$ users) loss. However their results assume they can solve a computationally hard mixed integer program optimally.
Preference Learning.
Given a fixed set of items, Tatli et al. (2022) focus on learning preferences distributions of users. Canal et al. (2022) extend this to the setting where the user losses are given by an identical metric, and they learn both preferences and the metric efficiently. Our setting focuses on the design of services rather than learning preferences over a fixed set of services, and we also allow each user to potentially have a different loss geometry; e.g., Mahalanobis distances with different matrices (where the sublevel sets of the loss are ellipsoids with different principal axes and orientations).
7 Discussion
It would be interesting to study how strategic user behavior affects the data acquisition and services’ initialization process. For example, users may report losses that are higher than their true losses in the hopes of being selected by the provider. This in turn can lead to a biasing of the service towards these particular untruthful users. It an interesting line of questioning centers on the design of algorithm modifications or the design explicit incentives to help avoid this, for example, by encouraging truthful reporting and also voluntary participation in the case that the designed mechanisms (algorithmic or incentive based) lead to users ‘opting out’ of all services.
On a related note, seemingly benign non-strategic behavior can also lead to degradation in performance. Specifically, there are many open questions relating to quantifying the robustness of the proposed initialization algorithm to noisy observations, perturbations, or outliers in the finite sample case when the feature distribution is heavy-tailed.
Finally, it is interesting to study the improvement in the total loss not only after initialization as was done in this paper, but also after running iterative learning dynamics, alternating service updates and users’ choice updates. It is known that the total loss will not increase and the iterations converge to a local minimum of the total loss, but further quantifying this improvement, as well as other properties of the iterative updates (including whether they can be carried out in our low-information setting) is left for future work.
@@ Acknowledgements
Dean, Fazel, Morgenstern and Ratliff were supported by NSF CCF-AF 2312774/2312775. Additionally, Morgenstern’s work was supported by NSF CCF 2045402. Dean’s research was supported by a gift from Wayfair, a LinkedIn research award, and NSF OAC 2311521. Ratliff’s research was supported by NSF CNS 1844729 and Office of Naval Research YIP Award N000142012571. Fazel was supported in part by awards NSF TRIPODS II 2023166, CCF 2007036, CCF 2212261.
Appendix A Mappings satisfying Assumptions 1 and 2
@@
In this appendix, we provide more details on examples that satisfy Assumptions 1 and 2. For the squared error loss for linear predictors, we refer the reader to Lemma 5 for a derivation of $c_{ij}$. The remainder of the example classes are detailed below.
A.1 Huber Loss
@@
A.2 The normalized cosine distance
@@
Assumption 2 is satisfied with $c_{ij}=\frac{1}{2}$ by using triangle inequality followed by power mean inequality.
A.3 Mahalanobis distance
@@
where $\Sigma_{i}$ is full rank and $c_{ij}=\min\{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j}),\frac{1}{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j})}\}$. The derivation is similar to proof of Lemma 5.
Appendix B Proof of Theorem 1
@@
The proof idea from here on is to show that there exists an approximation ratio $K_{\rm OPT}$ for the covered clusters, and is shown in Lemma 1 and 2.
Lemma 1.
@@
Using Assumption 2.(ii) with $\theta=\mathcal{J}(\mathcal{B})$, we have that
@@
By Assumption 2.$(i)$, for any $\theta\in\mathbb{R}^{d}$ we have that
@@
Note that $\sum_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\mathcal{L}_{i}(\phi_{j},\phi_{i})$ is essentially the loss on choosing users from $\mathcal{B}$ uniformly randomly and centering the service on the chosen user. Plugging in the expression in (5) from the proof of Lemma 1, we have that
@@
Replacing the factor 8 in Lemma 3.3 from (Arthur and Vassilvitskii, 2007) by $K_{\rm OPT}$ gives us the desired result.
With the preceding technical lemmas, we are now ready to prove the main theorem which we restate for convenience. See 1
@@
Consider $t=1$, and $\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})$, the cluster in which the first chosen user belongs. Applying Lemma 3 with $v=u=k-1$ and the fact that $\mathcal{B}$ is the only covered cluster, we have that
@@
by Lemma 1. Moreover, we have that
@@
Appendix C Fair Initialization
See 2
Proof.
For any $\Theta\subset\mathbb{R}^{d}$, we have that
$\Phi(\Theta,\mathcal{A})$ $=\max_{i\in[m]}\frac{\mathcal{L}(\Theta,\mathcal{A}_{i})}{|\mathcal{A}_{i}|}\leq\frac{1}{\min_{i\in[m]}|\mathcal{A}_{i}|}\sum_{i\in[m]}\mathcal{L}(\Theta,\mathcal{A}_{i})=\frac{1}{\min_{i\in[m]}|\mathcal{A}_{i}|}\mathcal{L}(\Theta,[n]).$
This implies for $\Theta_{k}$ the output of Algorithm 1, we have that
$\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]\leq\frac{1}{\min_{i\in[m]}|\mathcal{A}_{i}|}\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[n])].$
Hence, we have that
$\mathcal{L}(\Theta_{\rm fair},[n])$ $=\sum_{i\in[m]}\mathcal{L}(\Theta_{\rm fair},\mathcal{A}_{i})\leq\max_{i\in[m]}|\mathcal{A}_{i}|\sum_{i\in[m]}\max_{i\in[m]}\frac{\mathcal{L}(\Theta_{\rm fair},\mathcal{A}_{i})}{|\mathcal{A}_{i}|}=\max_{i\in[m]}|\mathcal{A}_{i}|m\cdot\Phi(\Theta_{\rm fair},\mathcal{A}).$
Now, since $\Theta_{\rm OPT}$ is the minimizer of $\mathcal{L}(\cdot,[n])$ and $\Theta_{\rm fair}$ the minimizer of $\Phi(\cdot,\mathcal{A})$, we have that
$\mathcal{L}(\Theta_{\rm OPT},[n])$ $\leq\max_{i\in[m]}|\mathcal{A}_{i}|m\cdot\Phi(\Theta_{\rm fair},\mathcal{A})$
$\leq\max_{i\in[m]}|\mathcal{A}_{i}|m\cdot\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]$
$\leq\frac{\max_{i\in[m]}|\mathcal{A}_{i}|}{\min_{i\in[m]}|\mathcal{A}_{i}|}m\cdot\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[n])].$
Using Theorem 1 for the guarantee on $\Theta_{k}$, we conclude the proof. ∎
C.1 What if User Demographic Identities are known?

### v2

@@ 5 Experiments
where $r_{i}$ is the true user ratings and $\Omega_{i}^{\rm train}$ is the list of movies rated by the users in subpopulation $i$. Since $r_{i}$ is a sparse matrix we consider the prediction error only on the movies they rated, i.e. $\Omega_{i}^{\rm train}$. User loss / dissatisfaction for a recommendation model, parameterized by $\theta\in\mathbb{R}^{d\times m}$, is captured by the excess error, namely
$\mathcal{L}_{i}(\theta,\phi_{i})=\|(U_{i}\theta-r_{i})_{\Omega_{i}^{\rm train}}\|_{2}^{2}-\|(U_{i}\phi_{i}-r_{i})_{\Omega_{i}^{\rm train}}\|_{2}^{2}.$
This value typically indicates how unhappy users are with the suggested movies with respect to their preferred movies. The provider initially doesn’t know the user ratings. At time step $t\leq k$, suppose the provider offers a set of recommendation models $\Theta_{t-1}$. Users choose the service with the best recommendations and the provider observes the losses across different subpopulations $\{\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[N]}$. Then the provider selects a subpopulation $l\in[N]$ to estimate $\phi_{l}$ via (5). This selection can be done via our proposed method or one of the baseline strategies discussed below. The provider then centers its next model on $\phi_{l}$ and updates the list of offered models $\Theta_{t}=\Theta_{t-1}\cup\phi_{l}$. In this process the provider only observes movie ratings of the users in the $k$ selected subpopulations. Once the services are initialized we evaluate the performance on the movies rated in the test set denoted by $\{\Omega_{1}^{\rm test},\ldots,\Omega_{N}^{\rm test}\}$.
Baselines. Both our tasks iterate through the steps of observing User Behavior, User Selection to gather data, designing New Service to update set of offered services. Through our experiments we wish to empirically evaluate different User Selection strategies with respect to AcQUIre (line 6 in Algorithm 1). The different user selection strategies result in the following baselines: (i) Random: $P(l=i)=1/n$, (ii) Greedy: $l=\operatorname*{\mathrm{argmax}}_{i\in[n]}\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})$, (iii) Epsilon Greedy: $l=\operatorname*{\mathrm{argmax}}_{i\in[n]}(\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})+\epsilon_{i})$ where $\epsilon_{1},\ldots,\epsilon_{n}$ denotes zero mean i.i.d. noise. Given that the Census Dataset comprises various racial demographic groups of varying sizes (with the smallest group being ten times smaller than the largest group), and considering our interest in the fairness objective (4), we explore incorporating these size imbalances into our algorithms. Consequently, we introduce three additional baselines, wherein the selection criteria are scaled by the corresponding group sizes: (iv) Balanced Random, (v) Balanced Greedy, and (vi) Balanced Epsilon Greedy. We benchmark them against Fair AcQUIre (Algorithm 2) which has guarantees as as stated in Theorem 3.2.
Evaluation: Each algorithm is run for 500 initialization seeds, the averages are reported in Figure 1.
Runtimes: We compare the runtimes of AcQUIre with the baselines, and study the affect of the number of users $(N)$ and number of services $(K)$ in Figure 2. We find the runtimes of AcQUIre to be of the similar order of magnitude of random initialization, meanwhile performing much better than random, and the much slower greedy and epislon greedy initialization schemes.
Impact of Initialization: Once a set of services are initialized, with more user interactions, the provider updates the services on new data to improve the quality (indicated by the reduction in total loss). To evaluate the importance of initialization, we conducted experiments using two different optimization algorithms: (i) Generalized k-means: The services are iteratively updated by training each service on the current group of subpopulations selecting it. After updating the service parameters, the subpopulations reselect their best service. This process repeats until convergence. (ii) Multiplicative weights update [12]: Similar to k-means, but each subpopulation can have users choosing different services simultaneously. Both generalized k-means and the multiplicative weights update guarantee that the total loss reduces over time [12].
In our experiments, we initialize a set of services using AcQUIre and other baseline methods, then let both optimization algorithms run until convergence. We plot the total loss versus the number of iterations (Figure 3). Our results demonstrate that AcQUIre leads to: (1) faster convergence, and (2) lower final loss (initializing with AcQUIre converges to lower losses; other initialization schemes are prone to being stuck in suboptimal local minima). These findings highlight the significance of a robust initialization strategy. By starting with a better initial configuration, the optimization algorithms can more effectively reach higher quality solutions.
Figure 3: We study the importance of initialization in both the convergence rate and quality of converged solution of optimization algorithms. We find AcQUIre converges both faster and to a lower total loss across optimization methods (kmeans and multiplicative weights) as well as datasets.
6 Conclusion
We study the problem of initializing services for a provider catering to a user base with diverse preferences. We address the challenges of unknown user preferences, only bandit (zeroth-order) feedback from the losses, and the non-convexity of the optimization problem, by proposing an algorithm that designs services by adaptively querying data from a small set of users. We also consider the fairness aspect of such design in human centric applications. Our proposed algorithm has theoretical guarantees on both the average and fair loss objectives. Directions for future work include quantifying the robustness of the proposed initialization algorithm to noisy observations, perturbations, or outliers in the finite sample case when the feature distribution is heavy-tailed.
@@ Acknowledgements
Dean, Fazel, Morgenstern and Ratliff were supported in part by NSF CCF-AF 2312774/2312775. Additionally, Morgenstern’s work was supported by NSF CCF 2045402. Dean’s research was supported by a gift from Wayfair, a LinkedIn research award, and NSF OAC 2311521. Ratliff’s research was supported by NSF CNS 1844729 and Office of Naval Research YIP Award N000142012571. Fazel was supported in part by awards NSF TRIPODS II 2023166, CCF 2007036, CCF 2212261.
Appendix A More Discussion on Related Works
Retention. User retention in machine learning systems is closely related to the decision dynamics between the provider and users studied in the single service setting by Hashimoto et al. [20], Zhang et al. [41] and multiple services setting by Dean et al. [12], Ginart et al. [18]. In settings with multiple sub-populations of users of different types, the question of retention has been explored in parallel with the issue of fairness. These works typically focus on the stability of the dynamics at equilibrium. However, the final outcome is heavily influenced by the initial data the provider has and the initial configuration (partitioning) of the users across the offered services. Our work addresses the impact of initialization on the final outcome with theoretical guarantees.
Mixture of Experts. This model specialization is also explored in the mixture of experts literature (see, e.g., [31]), which uses multiple ‘expert’ models to enhance accuracy and robustness, assigning inputs to the most suitable expert based on their features through a gating mechanism.
Clustering. [29] improved the analysis of [5] to show an approximation ratio $5(2+\log k)$ and showed a family of instances with $5\log k$ approximation ratio, thus showing $k$-means++ is tight. Recent works in clustering [24, 2, 26] provide a constant approximation ratio, and although they are polynomial time algorithms, these methods are data inefficient and rely on knowing the points a priori.
Facility Location Problem. Our algorithm has some resemblances with the facility location problem where there are $n$ users, and a provider can set up at most $k$ facilities at one of $m$ candidate locations (also known as the $k$-medoids objective [4]). One key difference is the provider can choose from $m$ predecided locations, compared to our setting where the optimization space is infinite. However our algorithm initializes only at $k$ of the $n$ user preferences, hence it can be viewed as the candidate locations simply being the user preferences and thus $m=n$. A typical greedy algorithm for the $k$-medoids objective proceeds by evaluating the marginal decrement in total loss for all possible candidates and selects the candidate with maximal loss reduction. Thus (i) the algorithm would need to know all $\{\phi_{1},\ldots,\phi_{n}\}$ apriori, and (ii) deploy $n$ services to obtain all $n^{2}$ function evaluations $\mathcal{L}_{j}(\phi_{i},\phi_{j}),\;\forall i,j\in[n]$. This is infeasible in applications like online recommendation systems where $n$ is very large and typically a provider has the capacity to deploy only $k\ll n$ services. Hence a $k$-medoids like objective will not be reasonable in our setup with incomplete information.
[6] studied the case where the provider had prior access to only $N<n$ users’ utility functions before deploying services and the goal was to minimize the worst case total (over all $n$ users) loss. However their results assume they can solve a computationally hard mixed integer program optimally.
Preference Learning. Given a fixed set of items or services, [38] focus on learning preference distributions of users. [8] extend this to the setting where the user losses are given by an identical metric, and they learn both preferences and the metric efficiently. Our setting focuses on the design of services rather than learning preferences over a fixed set of services, and we also allow each user to have a different loss function and loss geometry.
Appendix B Motivating Example
We conceptualize ‘services’ broadly, encompassing both sets of independent learners—such as various service providers collaborating on initialization—and single learners with multiple models, like companies with diverse platforms or multi-model servers in federated learning settings. This initialization process is further relevant for social planners aiming to facilitate the coordination of service initialization.
Netflix Example. All details are borrowed from the actual working of Netflix (refer https://recoai.net/netflix-recommendation-system-how-it-works/ for a detailed description and more references). The Netflix homepage displays several rows of suggestions. Each row is a collection of movies and the rows are arranged from top to bottom in decreasing order of likelihood that the user will pick movies to watch. The user has a complete choice of which row to select a movie to watch from. The different rows are generated by different underlying recommendation models parametrized by ${\theta_{1},\ldots,\theta_{K}}$. Our model abstracts these rows out as services. Netflix maintains $K=1300$ such different recommendation models. Users’ decision of watching a movie from the available rows of recommendations on the Netflix homepage informs the provider (Netflix) which of the rows (parameterized recommendation model) the user prefers most. This let’s them finetune their recommendation models.
Appendix C Mappings satisfying Assumptions 2.1 and 2.2
@@ Appendix B Motivating Example
In this appendix, we provide more details on examples that satisfy Assumptions 2.1 and 2.2. For the squared error loss for linear predictors, we refer the reader to Lemma F.1 for a derivation of $c_{ij}$. The remainder of the example classes are detailed below.
C.1 Huber Loss
@@
C.2 The normalized cosine distance
@@
Assumption 2.2 is satisfied with $c_{ij}=\frac{1}{2}$ by using triangle inequality followed by power mean inequality.
C.3 Mahalanobis distance
@@
where $\Sigma_{i}$ is full rank and $c_{ij}=\min\{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j}),\frac{1}{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j})}\}$. The derivation is similar to proof of Lemma F.1.
Appendix D Proof of Theorem 3.1
@@
The proof idea from here on is to show that there exists an approximation ratio $K_{\rm OPT}$ for the covered clusters, and is shown in Lemma D.2 and D.3.
Lemma D.2.
@@
Using Assumption 2.2.(ii) with $\theta=\mathcal{J}(\mathcal{B})$, we have that
@@
By Assumption 2.2.$(i)$, for any $\theta\in\mathbb{R}^{d}$ we have that
@@
Note that $\sum_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\mathcal{L}_{i}(\phi_{j},\phi_{i})$ is essentially the loss on choosing users from $\mathcal{B}$ uniformly randomly and centering the service on the chosen user. Plugging in the expression in (7) from the proof of Lemma D.2, we have that
@@
Replacing the factor 8 in Lemma 3.3 from [5] by $K_{\rm OPT}$ gives us the desired result.
With the preceding technical lemmas, we are now ready to prove the main theorem which we restate for convenience. See 3.1
@@
Consider $t=1$, and $\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})$, the cluster in which the first chosen user belongs. Applying Lemma D.4 with $v=u=k-1$ and the fact that $\mathcal{B}$ is the only covered cluster, we have that
@@
by Lemma D.2. Moreover, we have that
@@
Appendix E Fair Initialization